### v1

@@ 4   Evaluation Science
Disparities in true positive rates measure differences in receiving positive classifier decisions $\hat{Y}$, conditional on the underlying outcome being positive, $Y=1$, such as not defaulting on a loan (in lending prediction) or re-appearing for trial (in RAI prediction). One way to define an analogous interventional true positive rate is to posit that an individual should have received a positive decision $D=1$ (analogous to a positive classification label) if in fact they would have improved if they received the intervention. Predictive fair machine learning posits the true positive label $Y=1$ as sufficient information to “merit” a positive classification decision $\hat{Y}=1$. An interventional analogue is to posit that those who strictly benefit, i.e. respond under the intervention $Y(1)>Y(0)$ (Kallus and Zhou, 2019), ought to receive the intervention, $D=1$. Imai and Jiang (2023) also develops fairness criteria based on these principal strata. A true positive rate disparity across groups can be defined as follows.
@@
We first discuss quasi-experimental methods such as event studies based on pre-post comparisons prior to or after the deployment of ADS, or other quasi-experimental methods applicable when micro-data is available but decisions $D$ are a known function of $R$.
@@
One way to model the causal impacts of ADS is to view the introduction of the ADS itself as a policy intervention or “program”, considering the impacts on relevant downstream outcomes prior vs post the introduction of an ADS tool into practice. Often these studies use event-study pre-post designs on outcomes before and after the introduction of the ADS. Usually event study methods like difference-in-differences and synthetic control are particularly vulnerable to underlying statistical model assumptions, since they usually lack experimental randomness.
@@
Case Study 2 (Case Study Two: Observational quasi-experimental evaluation).
Albright (2019) assesses observational case-level data sourced from the Kentucky Administrative Office of the Courts, before and after a certain Kentucky policy change on June 8, 2011- House Bill 463 (HB463). The policy requires judges to more actively factor in the assessment score level (high, moderate and low), as calculated by the Kentucky Pretrial Risk Assessment (KPRA), in their decision-making. They were requested to release all low and moderate risk defendants with a non-financial bond, unless actively providing justification for their deviance.
The author first analyzes trends in the data from 2009-2013, identifying a discontinuity in the percentage of non-financial bonds after the policy implementation. They conclude that although the explicit consideration of the risk score via HB463 overall leads to more leniency in judge decision-making, the scores have little to no effect on high risk defendant decisions of any race and Black defendants in general. A difference-in-difference analysis reveals that pre-HB463, differences between Black and White defendants is negligible. However, post-HB463, white defendants receive a significantly greater amount of leniency, experiencing 30% and 62% more of the short-term gains in non-financial bonds than their Black counter-parts in low and moderate risk score settings. Even when adjusting for time-dependent judge-fixed effectsBlack defendants still experience a 25% lower increase in non-financial bond issuance.
Event-studies are often used to study the impacts of non-randomized broad policy interventions. Broad policies can have complex relationships with specific ADS and decision-making patterns. (Zhou et al., 2021) conduct an event-study analysis of initial bail reform efforts in New York; but the impacts of bail reform on the different distribution of released defendants are complex.
Such studies based on before-after contrasts of outcomes can measure aggregate impacts of introducing ADS at all. However, this type of event-study analysis typically doesn’t consider the per-individual information of how the ADS predictions specifically impacted decisions.
Evaluating ADS requires understanding both their effectiveness and the mechanisms behind their impact. An important question when evaluating ADS is whether using them affects decisions and outcomes. However, measuring such impacts is only one part of the broader question of why they work. A decision-support tool that perfectly predicts the treatment effects for individuals may lead to no change in outcomes if decision-makers ignore it. On the other hand, even weak predictors could hypothetically improve outcomes by increasing oversight or emphasizing consistency.
Disentangling different mechanisms through which ADS operate can pin down vastly different implications for next steps. Should an organization investigate what informs front-line decision-making in more detail, or should they develop more effective programs/interventions to allocate? Disaggregating evaluations based on individual-level micro-data decisions and outcomes can provide evidence as to mechanisms. We next discuss different evaluation frameworks for doing so.
@@
Such a definition can be derived within a decision-theoretic framework similar to that discussed above, as the rate at which an idealized decision-maker with access to the human’s independent predictions and those of a predictive model relies on the model. Consider a rational Bayesian agent who sees the independent human judgment and the model prediction as part of the “signal” available at the time of the decision (in addition to features $X$, model explanations, and other provided information), and must choose which to go with as the final decision. This agent knows the prior probability that the model is correct over the distribution of interest, and has oracle access to the joint distribution over signals and states. For any decision instance they encounter, they use their knowledge of the joint distribution to update their beliefs to Bayesian posterior beliefs, and choose the score-maximizing action under the utility function associated with the decision problem (Guo et al., 2024). The performance of this agent benchmarks the maximum attainable performance of a human deciding between their own and the model’s prediction. (McLaughlin and Spiess, 2022) study a rational decision-making perspective with preferences, while (McLaughlin and Spiess, 2024) leverages the potential outcomes framework.
@@
Beyond assessing the potential for complementary performance and benchmarking how close a human-AI pair comes to the upper bound on performance, we are often interested in how and how well agents (human and artificial) exploit available information. A longstanding, robust finding in decision studies is that when statistical prediction is evaluated head-to-head against clinical (i.e., human prediction), the statistical model typically exceeds or matches the human experts’ performance Ægisdóttir et al. (2006); Grove et al. (2000); Meehl (1954). This finding goes against common intuitions that human decision-makers can possess valuable private information that is hard to encode in statistical models, such as through interacting directly with the decision subject. Recent works explore the presence and value of information uniquely available to humans or other agents. Alur et al. (2024) develops a statistical hypothesis testing to test whether human predictions are statistically independent from the outcomes after conditioning on the features, where a rejection implies that the human decision maker may have additional expertise beyond the information available to a model (algorithm). Schechtman et al. (2025) extends this work from prediction-only settings to interventions, presenting a causal graphical framework for detecting expertly targeted action and showed the value of mixed-methods analyses of human discretion in the context of AI-supported college advising. Guo et al. (2025) present a value of information theoretic framework for measuring the marginal value of available information (such as instance features or another agents’ predictions) for decisions.
@@
A popular tool is to use ideas from regression discontinuity design to leverage the deployment of a predictive algorithm for the purposes of evaluating the causal impacts of that decision. [^19] A closely related design is the tie-breaking design: capacity constraints that might result in randomization at the final threshold of eligibility. For example, Smith (2002) is an early use of this design to evaluate the impacts of “profiling” unemployment insurance seekers by their risk of lengthy unemployment spells and targeting them for mandatory re-employment services. Although typical regression discontinuity designs consider a cutoff on a one-dimensional “running-variable” (i.e. individual receives treatment when the “running-variable”, a score, is greater than some value), machine-learning based algorithms generate high-dimensional decision boundaries. RDDs work by contending that individuals right “above” or “below” treatment have quite similar characteristics, and therefore differences in outcomes can be attributable only to the difference in treatment (rather than differences in characteristics). The typical setup for an RDD posits that an individual crossing the threshold exactly receives treatment, which corresponds to assuming that algorithmic classifiers or risk scores determine decisions; however it can require additional data collection to record prior predicted risk scores (Perdomo et al., 2023; Gallo et al., 2024; Cowgill, 2018).
It is rarely the case that predictive risk models themselves impact final outcomes since they often only inform decisions, as in the case of algorithmic decision support. When they do not determine decisions, there is an analogous notion of “fuzzy RDDs”. Although RDDs are generally compelling, causal effects are only defined locally for individuals with covariates right at the threshold.
@@
So far, we have discussed a whole range of strategies to use in observational data. In causal inference, the strongest evidence is typically considered to be a randomized controlled trial (RCT). Since the impacts of ADS stem from how the predictions influence actions, randomized experiments and in-deployment evaluations provide critical evidence as to what happens in the absence or presence of ADS, grounding discussions of potential unintended consequences in empirical observation. Introducing random variation in whether a decision instance receives ADS information may be the ideal RCT.
Observational studies also inform or complement randomized evaluations. Methods for combining observational and experimental methods via data fusion (Bareinboim and Pearl, 2016; Colnet et al., 2024) can achieve the best of both worlds between the plentiful availability (but weaker identification) or observational data vs. the data scarcity (but stronger identification) of randomized data.
Designing and implementing experimental evaluations for ADS in view of these complexities is an open area of research. The next case study outlines a RCT of provision of ADS (the PSA) and an example of interventional evaluation: metrics derived on the principal strata of whether someone decides to give the decision or not based on whether they see the PSA or not (and analogously for outcomes).
@@
During a 30-month assignment period, spanning 2017 until 2019, judges in Dane County, Wisconsin were either given or not given a pretrial Public Safety Assessment (PSA) score for a given case. Although the PSA was calculated for every case, randomization determined whether the judge saw the score  (Imai et al., 2023). Given these scores, the judge made the decision $D_{i,k}$ to either enforce a signature bond, a small ($<$\$1000) or large case bail ($>$\$1000). Information about judge decisions, and defendant outcomes, $Y_{i,k}$ (ie. failure to appear (FTA), new criminal activity (NCA) and new violent criminal activity (NVCA)) are tracked for a period of two years after randomization.
@@
So far, we have emphasized singular choices of outcomes along which to assess whether ADS leads to improved decision-making. However, what constitutes a “better” decision or decision-making process is often underspecified and contested.
As we have observed throughout, ADS address complex social issues without clear solutions, affecting stakeholders with diverse interests and commitments. Many institutional decisions, especially in high-stakes public-sector contexts, involve multiple, sometimes conflicting, higher-level objectives (Green and Chen, 2021; Zacka, 2017). For instance, college enrollment algorithms that rank applicants for admissions or financial assistance might be evaluated according to their original purpose – often, optimizing yield – but good college admissions decision-making may serve additional goals valued by stakeholders, such as ensuring student academic success or enhancing campus diversity (Engler, 2021).
@@
In the field of program evaluation, there is a major emphasis on the development of performance standards. The proper formulation of evaluations in the context of a machine learning task, including the choice of outcomes to measure, presents a similar challenge. In particular, in many legal product safety regimes, the clear articulation of system expectations is essential in setting up robust audits and performance verification schemes (Raji et al., 2022a, b).
  • Internal Standards There is a long-standing tradition across various fields of internal auditing and pre-deployment testing Raji et al. (2020). These are internally set principle documents and guidelines that articulate the expectations set by organizational leadership, which can then inform local procurement and model testing.
  • Proprietary Standards Several standards are developed and sold by standards-making bodies, such as International Organization for Standardization (ISO), or the Institute of Electrical and Electronics Engineers (IEEE). Guiding documents and templates such as NIST’s Risk Management Framework (RMF), Microsoft’s AI Fairness Checklist or Google’s Model Cards Mitchell et al. (2019), which is broadly disseminated as a type of evaluation tool across a range of organizations Ojewale et al. (2024).
  • Government Standards In many cases, the expectations of the system are set by law. This can include, for example, fairness criteria set by legislative anti-discrimination statues  (Xiang, 2020; Wachter et al., 2020; Xiang and Raji, 2019), or product safety standards of performance set by rule-making enforcement agencies  Raji et al. (2022a). Although more rigid, these expectations also tend to be more vague and open to interpretation, allowing for some flexibility in how adherence to the standard is contextualized within a broader accountability process. For instance, the fourth-fifths heuristic for employment discrimination is only the first step in a multi-step process of legal accountability  Raghavan and Kim (2024).
@@
A third critical challenge in evaluating ADS is that measurement itself can influence or distort the behaviors and processes being assessed. This dynamic is captured by Goodhart’s law: “when a measure becomes a target, it ceases to be a good measure” (Goodhart, 1984). Once actors – whether individuals, organizations, or entire institutional systems – know they are being assessed on a particular metric, they may optimize for that metric in ways that distort its intended purpose. Scholars from across disciplines have documented the challenge of differentiating between genuine improvements in performance and strategic gaming of metrics without actual progress on the underlying dimension being measured (Hardt et al., 2016; Kleinberg and Raghavan, 2019; Laufer et al., 2023b). This issue is particularly pronounced when evaluative measures intersect with professional incentives or performance measurement systems. Qualitative evidence from Saxena et al. (2021a) also highlights that algorithms used in the child welfare risk assessment ecosystem often lead to “gaming” behaviors, such as caseworkers manipulating the inputs to achieve the desired outcome or manage conflict, as that is the only way for them to regain agency within a system that limits their discretion.
@@ 5   Implementation Science
In this section, we describe the challenge of “implementation design”, which is effectively oriented around the integration of context - ranging from the narrow context of human-AI interactions to a broader institutional and societal context - in deployment practices of machine learning systems.
@@
In some settings the human and the model have access to potentially differing information, and human control over the final decision is not required (quadrant 3). Here, the question becomes how to best distribute test instances to the two types of agents. One approach is rejection learning (Cortes et al., 2016), where the model’s confidence in its prediction is used to put aside instances that the model has high uncertainty over, which are given to a human. However, better performance can be gained by a learning to defer framework that jointly learns a predictive model and a rejector that determines whether the classifier or the human should predict (Mozannar and Sontag, 2020; Mozannar et al., 2023), as this approach takes into account the human’s expected performance as well.
When the human and the model have access to potentially differing information, but the human is expected to retain final control over the decision (quadrant 4), the challenge becomes how to best inform the human decision-maker about their expected performance versus that of the prediction model. Using feedback from realized outcomes to extract information on human versus model performance over the feature space could inform probabilistic predictions on the current instance. Methods for quantifying prediction uncertainty, such as conformal prediction or posthoc scaling of model-estimated confidence like softmax, must be applied with care when the human may have access to unique information Hullman et al. (2025c). Calibration of methods should take into account both agents’ chances of being correct to prevent misleading the decision-maker . One way to achieve is this through multi-calibrated uncertainty (Corvelo Benz and Rodriguez, 2024): calibrating probabilistic predictions over the model’s information as well as the human’s confidence function. Model explanations can similarly be useful to decision-makers with unique information over the model if they allow the decision-maker to compare their own internal model of the process to that implied by the model prediction.
@@
Another critical aspect of individual context is that of designing the intervention to be executed in response to the ADS deployment. In order for the ADS integration within a broader system to be effective, this intervention must also be designed and optimized to be the most meaningful and effective.
For example, designing interventions in response to ADS within education requires more than simply identifying at-risk students; it demands thoughtful, evidence-based interventional strategies. Liu et al. (2023b) challenged the assumption that surfacing risk scores is inherently useful, arguing that ‘naive’ interventions such as showing risk scores to students and schools without providing additional resources and support can hurt more than help. McConvey et al. (2023) found that interventions in response to ADS in higher education took various forms, including automated email notifications, supplemental coursework, and advising meetings, but none showed consistent empirical evidence of improving student outcomes. They highlight the potential of iterative, participatory design approaches that involve learning strategists and other stakeholders in developing theoretically grounded interventions alongside ADS model design.
@@
Beyond the “bureaucratic counter-factual” as a baseline institutional policy, an ADS deployment is often even more complex in practice. At any point, various independent actors can intervene and shape outcomes at various stages of the decision-making process, and make use of one or more ADS tools to do so. The result is an institutional policy change involving multiple steps and multiple stakeholders, each of which can, in some way, impact institutional outcomes (See Figure  8 for such a journey map illustrating this complexity).
@@
In order to properly dig into this question and expose in detail the complexity of how the deployment context can shape system outcomes, we start with a simple example scenario - “ What is the difference between predicting risk in (A) criminal recidivism and (B) hospital re-admission?” These prediction tasks (as well as similar tasks such as e.g. child-welfare alert algorithms (Saxena et al. (2023)), have an almost identical problem formulation – input characteristics of the model subject predict the probability of a negative outcome (i.e. the inappropriate release of the model subject, requiring re-entry into the institutional system) and thus inform interventions to minimize that negative outcome in practice. Although the prediction task of clinical re-admission and criminal recidivism seem on their face to have similar properties (ie. how likely is the release/discharge to fail and the model subject to return?), we can demonstrate how several external contextual factors, conditions of deployment and integration practices can lead the prediction models to have vastly different degrees of impact in-situ.
@@
As an example, the “journey maps” for the implementation of the Duke Sepsis Watch ADS system are incredibly complex  (Boag et al., 2024) and illustrates the wide variety of involved stakeholders. The journey map spans from the executive hospital management who approve funding, to the nurse practitioners who develop and train on-the-ground users, to the regulatory compliance teams who vet the model’s suitability for deployment given legal and institutional restrictions. The fact is, like any product or policy change, well-implemented prediction-based interventions involve a detailed change management plan to ensure that the set of expectations under a new regime are met, and unexpected challenges are directly addressed.
Bureaucratic Counter-Factuals & Policy Decision-Making Pipelines
Most predictions are made in the context of an existing policy implementation. This means that the outcome of a policy before the prediction is the result of a sequence of decisions, each of which informs the probability of another set of choices. The judges, $J_{k,i}$ making such decisions, $D_{i}$ are at times distinct from the actors $A_{l,i}$ acting on an intervention $I_{i}$. These actors are often distinct from the decision subject $S_{m},i$ and the creator, procurer or developer of the model prediction $C_{j},i$. In reality, there are multiple stakeholders involved at different decision and action states, and fully mapping out the default bureaucratic policy with and without the introduction of a prediction system can be quite the complex endeavor.
In addition to this, rarely is there just a single model prediction $\hat{Y}$ introduced at a single decision-making point $D_{i}$. Rather, most deployments manifest as multi-prediction pipelines – that is, multiple models used at different stages of a complex decision-making process, interacting with distinct stakeholders and other models as part of a systematic ADS pipeline (Saxena and Guha, 2024).
@@
We explore the complexity of how the deployment context can shape system outcomes by asking, “What is the difference between predicting risk in (A) criminal recidivism and (B) hospital re-admission?” The research literature presents these prediction tasks as being almost identical in problem formulation (as well as similar tasks such as e.g. child-welfare alert algorithms (Capatosto (2017), homeless service prioritization algorithms such as VI-SPDAT, etc).
In both of these cases, the prediction task can be characterized as the likelihood of an inappropriate release of the model subject, requiring re-entry into the institutional system, i.e. how likely is the release/discharge to fail and the model subject to return? However, although the prediction task of predicting clinical re-admission and criminal recidivism seem on their face to have similar properties, we can demonstrate how the differing context of the two scenarios leads to deployment and integration practices that drastically shape the degree to which the predictions shape system outcomes. Moreover, this problem is further complicated within specific domains. For instance, within child welfare systems, predicting risk of re-referral and risk of discharge often use similar data inputs and modeling strategies yet, qualitatively mean very different things to the decision-makers implementing interventions in-situ ((Saxena et al., 2023)).
@@
What are the stakeholders that impact the outcome?  [^21] Note that each of these stakeholders can preside over multiple cases $i$.For instance, $S_{m,i}$ can be a decision subject over time across multiple cases $i$.

### v2

@@ 4  Evaluation Science
Disparities in true positive rates $P(\hat{Y}=1\mid Y=1,A=a)$ measure differences in receiving positive classifier decisions $\hat{Y}$ conditional on having the underlying “positive label” such as not defaulting on a loan (in lending prediction) or re-appearing for trial (in RAI prediction). One way to define an analogous interventional true positive rate is to posit that an individual should have received a positive decision $D=1$ if in fact they would have improved if they received the intervention. The interventional analogue posits that those who strictly benefit, i.e. respond under the intervention $Y(1)>Y(0)$ (Kallus and Zhou, 2019), ought to receive the intervention, $D=1$. Imai and Jiang (2023) also develops fairness criteria based on these principal strata. A true positive rate disparity across groups can be defined as follows.
@@
One way to model the causal impacts of ADS is to view the introduction of the ADS itself as a policy intervention or “program”, considering the impacts on relevant downstream outcomes prior vs post the introduction of an ADS tool into practice. Often these studies use event-study pre-post designs on outcomes before and after the introduction of the ADS.
@@
Case Study 2 (Before-after comparison with panel data and quasi-experimental methods).
Albright (2019) assesses observational case-level data sourced from Kentucky before and after House Bill 463 (HB463) which required judges to more actively factor in the the Kentucky Pretrial Risk Assessment (KPRA)’s risk label (high, moderate and low) in their decision-making. They were requested to release all low and moderate risk defendants with a non-financial bond unless otherwise justified.
The author first analyzes temporal trends and identifies a discontinuity in the percentage of non-financial bonds after the policy implementation. They conclude that considering the ADS overall leads to more leniency in judge decision-making, the scores have little to no effect on decisions on high-risk defendant decisions of any race, including black. A difference-in-difference analysis reveals that pre-HB463, differences between Black and White defendants are negligible. However, post-HB463, white defendants receive a significantly greater amount of leniency, experiencing 30% and 62% more of the short-term gains in non-financial bonds than their Black counter-parts in low and moderate risk score settings. Even when adjusting for time-dependent judge-fixed effectsBlack defendants still experience a 25% lower increase in non-financial bond issuance.
Event-studies are often used to study the impacts of non-randomized broad policy interventions. Broad policies can have complex relationships with specific ADS and decision-making patterns.
Such studies based on before-after contrasts of outcomes can measure aggregate impacts of introducing ADS at all232323(Zhou et al., 2021) conduct an event-study analysis of initial bail reform efforts in New York; but the impacts of bail reform on the different distribution of released defendants are complex.. However, this type of event-study analysis typically doesn’t consider the per-individual information of how the ADS predictions specifically impacted decisions. Evaluating ADS requires understanding both the magnitude of their overall effectiveness, and disaggregated along the potential mechanisms behind their impact. Implications for next steps could be vastly different: should an organization investigate what informs front-line decision-making in more detail, or should they develop more effective programs/interventions to allocate? We next discuss different evaluation frameworks for disaggregating evaluations based on individual-level micro-data decisions and outcomes.
@@
Such a definition can be derived within a decision-theoretic framework similar to that discussed above, as the rate at which an idealized decision-maker with access to the human’s independent predictions and those of a predictive model relies on the model. Consider a rational Bayesian agent who sees the independent human judgment and the model prediction as part of the “signal” available at the time of the decision (in addition to features $X$, model explanations, and other provided information), and must choose which to go with as the final decision. The ideal agent knows the prior probability of model correctness and the joint distribution over signals and states: they can then conduct Bayesian posterior updating and take belief-optimal actions (Guo et al., 2024). The performance of this agent benchmarks the maximum attainable performance of a human deciding between their own and the model’s prediction. (McLaughlin and Spiess, 2022) study a rational decision-making perspective with preferences, while (McLaughlin and Spiess, 2024) leverages the potential outcomes framework.
@@
Beyond assessing the potential for complementary performance and benchmarking how close a human-AI pair comes to the upper bound on performance, we are often interested in how and how well agents (human and artificial) exploit available information. A longstanding, robust finding in decision studies is that when statistical prediction is evaluated head-to-head against clinical (i.e., human prediction), the statistical model typically exceeds or matches the human experts’ performance Ægisdóttir et al. (2006); Grove et al. (2000); Meehl (1954). This finding goes against common intuitions that human decision-makers can possess valuable private information that is hard to encode in statistical models, such as through interacting directly with the decision subject. Recent works explore the presence and value of information uniquely available to humans or other agents (Alur et al., 2024; Schechtman et al., 2025; Guo et al., 2025).
@@
A popular tool is to use ideas from regression discontinuity design to leverage the deployment of a predictive algorithm for the purposes of evaluating the causal impacts of that decision (Perdomo et al., 2023; Gallo et al., 2024; Cowgill, 2018). [^24] A closely related design is the tie-breaking design: capacity constraints that might result in randomization at the final threshold of eligibility. For example, Smith (2002) is an early use of this design to evaluate the impacts of “profiling” unemployment insurance seekers by their risk of lengthy unemployment spells and targeting them for mandatory re-employment services.
Since predictive risk models rarely themselves impact final outcomes, since they often only inform decisions, “fuzzy RDDs” are typically more appropriate. Although RDDs are generally compelling, causal effects are only defined locally for individuals with covariates right at the threshold.
@@
So far, we have discussed a whole range of strategies to use in observational data. In causal inference, the strongest evidence is typically considered to be a randomized controlled trial (RCT). Introducing random variation in whether a decision instance receives ADS information may be the ideal RCT.
Designing and implementing experimental evaluations for ADS in view of these complexities is an open area of research. The next case study outlines a RCT of provision of ADS (the PSA) and an example of interventional evaluation via principal stratification.
@@
During a 30-month assignment period, spanning 2017 until 2019, judges in Dane County, Wisconsin were either given or not given a pretrial Public Safety Assessment (PSA) score for a given case. Although the PSA was calculated for every case, randomization determined whether the judge saw the score  (Imai et al., 2023). Given these scores, the judge made the decision $D_{i,k}$ to either enforce a signature bond, a small (<<\$1000) or large case bail (>>\$1000). Information about judge decisions, and defendant outcomes, $Y_{i,k}$ (ie. failure to appear (FTA), new criminal activity (NCA) and new violent criminal activity (NVCA)) are tracked for a period of two years after randomization.
@@
So far, we have emphasized singular choices of outcomes along which to assess whether ADS leads to improved decision-making. However, what constitutes a “better” decision or decision-making process is often underspecified and contested. As we have observed throughout, ADS address complex social issues without clear solutions, affecting stakeholders with diverse interests and commitments. Many institutional decisions, especially in high-stakes public-sector contexts, involve multiple, sometimes conflicting, higher-level objectives (Green and Chen, 2021; Zacka, 2017). [^25] For instance, college enrollment algorithms that rank applicants for admissions or financial assistance might be evaluated according to their original purpose – often, optimizing yield – but good college admissions decision-making may serve additional goals valued by stakeholders, such as ensuring student academic success or enhancing campus diversity (Engler, 2021).
@@
In the field of program evaluation, there is a major emphasis on the development of performance standards. The proper formulation of evaluations in the context of a machine learning task, including the choice of outcomes to measure, presents a similar challenge. In particular, in many legal product safety regimes, the clear articulation of system expectations is essential in setting up robust audits and performance verification schemes (Raji et al., 2022a, b). This is also closely connected to performance management. Under the U.S. Government Performance and Results Act, for instance, agencies must define and report performance measures. Many agencies institute quality assurance programs – often utilizing random samples of cases – to measure and improve outcomes (Ames et al., 2020).
Internal standards are the starting point via internal auditing and pre-deployment testing Raji et al. (2020), articulating organization expectations to inform local procurement and model testing. Proprietary standards include those developed by standards-making bodies, such as International Organization for Standardization (ISO), or the Institute of Electrical and Electronics Engineers (IEEE), or other widely disseminated guiding documents such as NIST’s Risk Management Framework (RMF), Microsoft’s AI Fairness Checklist or Google’s Model Cards Mitchell et al. (2019), widely disseminated across a range of organizations Ojewale et al. (2024). In many cases, the expectations of the system are set by law in governmental standards. This can include, for example, fairness criteria set by legislative anti-discrimination statues  (Xiang, 2020; Wachter et al., 2020; Xiang and Raji, 2019), or product safety standards of performance set by rule-making enforcement agencies  Raji et al. (2022a).
@@
A third critical challenge in evaluating ADS is that measurement itself can influence or distort the behaviors and processes being assessed. This dynamic is captured by Goodhart’s law: “when a measure becomes a target, it ceases to be a good measure” (Goodhart, 1984). Once actors – whether individuals, organizations, or entire institutional systems – know they are being assessed on a particular metric, they may optimize for that metric in ways that distort its intended purpose. Scholars from across disciplines have documented the challenge of differentiating between genuine improvements in performance and strategic gaming of metrics without actual progress on the underlying dimension being measured (Hardt et al., 2016; Kleinberg and Raghavan, 2019; Laufer et al., 2023b). This issue is particularly pronounced when evaluative measures intersect with professional incentives or performance measurement systems. Qualitative evidence from Saxena et al. (2021a) also highlights that algorithms used in the child welfare risk assessment ecosystem often lead to “gaming” behaviors, such as caseworkers manipulating the inputs to achieve the desired outcome or manage conflict, as that is the only way for them to regain agency within a system that limits their discretion. [^26] Similarly, Ho et al. (2019) documents that the Board of Veterans Appeals lowered the standard of review to be able to report consistently high “accuracy” of adjudicated cases, posing a sharp tension between the goals of error correction and performance reporting.
@@ 5  Implementation Science
In this section, we describe the challenge of “implementation design”, which integrates context - ranging from the narrow context of human-AI interactions to a broader institutional and societal context - in deployment practices of machine learning systems.
@@
In some settings the human and the model have access to potentially differing information, and human control over the final decision is not required (quadrant 3). Here, the question becomes how to best distribute test instances to the two types of agents. One approach is rejection learning (Cortes et al., 2016) which puts aside instances with high predictive uncertainty. However, better performance can be gained by a learning to defer framework that jointly learns a predictive model and a rejector that determines whether the classifier or the human should predict (Mozannar and Sontag, 2020; Mozannar et al., 2023), as this approach takes into account the human’s expected performance as well.
When the human and the model have access to potentially differing information, but the human is expected to retain final control over the decision (quadrant 4), the challenge becomes how to best inform the human decision-maker about their expected performance versus that of the prediction model. Using feedback from realized outcomes to extract information on human versus model performance over the feature space could inform probabilistic predictions on the current instance. Methods for quantifying prediction uncertainty, such as conformal prediction or posthoc scaling of model-estimated confidence like softmax, must be applied with care when the human may have access to unique information Hullman et al. (2025c). Calibration of methods should take into account both agents’ chances of being correct to prevent misleading the decision-maker, for example with multi-calibrated uncertainty (Corvelo Benz and Rodriguez, 2024) over the model’s information as well as the human’s confidence function.
@@
Another critical part of individual context is the specific intervention design of what would be done in response to ADS information. Institutions must also consider designing interventions for effective ADS integrations.
For example, improving outcomes in education with ADS requires not only simply identifying at-risk students, but also thoughtful, evidence-based interventional strategies. Liu et al. (2023b) challenged the assumption that surfacing risk scores is inherently useful, arguing that ‘naive’ interventions such as showing risk scores to students and schools without providing additional resources and support can hurt more than help. McConvey et al. (2023) found that interventions in response to ADS in higher education took various forms, including automated email notifications, supplemental coursework, and advising meetings, but none showed consistent empirical evidence of improving student outcomes. They highlight the potential of iterative, participatory design approaches that involve learning strategists and other stakeholders in developing theoretically grounded interventions alongside ADS model design.
@@
Beyond the “bureaucratic counter-factual” as a baseline institutional policy, an ADS deployment is often even more complex in practice. ADS integrationss are an institutional policy change involving multiple steps and multiple stakeholders, each of which can, in some way, impact institutional outcomes (See Figure  8 for such a journey map illustrating this complexity).
@@
As an example, the “journey maps” for the implementation of the Duke Sepsis Watch ADS system are incredibly complex  (Boag et al., 2024) and illustrates the wide variety of involved stakeholders. The journey map spans from the executive hospital management who approve funding, to the nurse practitioners who develop and train on-the-ground users, to the regulatory compliance teams who vet the model’s suitability for deployment given legal and institutional restrictions. Well-implemented prediction-based interventions need a detailed change management plan to ensure meeting new expectations and addressing unexpected challenges.
Bureaucratic Counterfactuals & Policy Decision-Making Pipelines
In reality, the bureaucratic counterfactual we have been referring to in abstract comprises multiple process steps and stakeholders. Interventions made without predictions still require a sequence of decisions, each of which informs the probability of another set of choices. The judges, $J_{k,i}$ making such decisions, $D_{i}$ are at times distinct from the actors $A_{l,i}$ acting on an intervention $I_{i}$. These actors are often distinct from the decision subject $S_{m},i$ and the creator (procurer or developer) of the model prediction $C_{j},i$. Fully mapping out the bureaucratic counterfactual with or without ADS can be complex.
Further, most deployments manifest as multi-prediction pipelines, not single model predictions at a single decision. They require multiple models used at different stages of a complex decision-making process, interacting with distinct stakeholders and other models as part of a systematic ADS pipeline (Saxena and Guha, 2024).
@@
Considering a simple question highlights the complexity of how the deployment context can shape system outcomes: “What is the difference between predicting risk in (A) criminal recidivism and (B) hospital re-admission?” These prediction tasks (as well as similar tasks such as e.g. child-welfare alert algorithms (Saxena et al. (2023)), have an almost identical problem formulation – input characteristics of the model subject predict the probability of a negative outcome (i.e. the inappropriate release of the model subject, requiring re-entry into the institutional system) and thus inform interventions to minimize that negative outcome in practice. Although the prediction task of clinical re-admission and criminal recidivism seem on their face to have similar properties (ie. how likely is the release/discharge to fail and the model subject to return?), several conditions of deployment and integration practices can lead the prediction models to have vastly different degrees of impact in-situ.
Moreover, even within specific domains, prediction tasks that sound like similar prediction tasks yield vastly different decision-making processes. For instance, within child welfare systems, predicting risk of re-referral and risk of discharge often use similar data inputs and modeling strategies yet, qualitatively mean very different things to the decision-makers implementing interventions in-situ ((Saxena et al., 2023)).
@@
What are the stakeholders that impact the outcome?  [^27] Note that each of these stakeholders can preside over multiple cases ii.For instance, $S_{m,i}$ can be a decision subject over time across multiple cases ii.